# RaSeRec: Retrieval-Augmented Sequential Recommendation

### v2

@@ 4.1 Experimental Settings
Baselines. To verify the effectiveness of RaSeRec, we compare it with the following three groups of competitive models: (1) Non-sequential recommendation models (Non-SeRec) include PopRec and BPR-MF Rendle et al. (2009); (2) Vanilla sequential recommendation models (Vanilla SeRec) include GRU4Rec Hidasi et al. (2016), Caser Tang and Wang (2018), SASRec Kang and McAuley (2018), BERT4Rec Sun et al. (2019), and $\rm S^{3}Rec_{MIP}$ Zhou et al. (2020); (3) SSL-Augmented sequential recommendation models (SSL-Augmented SeRec) include CL4SRec Xie et al. (2022), CoSeRec Liu et al. (2021), ICLRec Chen et al. (2022b), DuoRec Qiu et al. (2022), and MCLRec Qin et al. (2023). Detailed descriptions of each baseline model can be seen in Appendix C.
@@
Implementation and Evaluation. We use Recbole Zhao et al. (2022) to implement baselines. For models with learnable item embedding, we set the hidden size $d=64$. For each baseline, we follow the suggested settings reported in their original paper to set hyper-parameters. More implementation details can be seen in Appendix E. Following the leave-one-out strategy, we hold out the last interacted item of each user sequence for testing, the second last item for validation, and all the earlier ones for training. We measure the recommendation performance via HR and NDCG (see Appendix D), where the cut-off position (@$N$) is set as 5 and 10.
4.2 Comparison with Baselines (RQ1)
@@
Table 2 and 8 summarize the performance of all baselines and RaSeRec w.r.t. @5 and @10, respectively. From the results, we observe that: (1) The performance of the Non-SeRec models is unsurprisingly worse than the vanilla SeRec models, indicating the necessity of modeling sequential patterns for capturing more accurate user preferences. On the other hand, SSL-Augmented SeRec models perform better than those without SSL, indicating the superiority of supplementing the recommendation task with self-supervised learning. (2) Benefiting from the retrieval-augmented mechanism, RaSeRec considerably outperforms all baseline models in 16 out of 18 cases with $p$-value less than 0.005, demonstrating the improvement of RaSeRec over the SOTA baselines is statistically significant. The reasons are mainly twofold: (i) By sustainably maintaining a dynamic memory bank, RaSeRec can quickly adapt to preference drifts. (ii) By retrieving memories, RaSeRec can explicitly recall long-tailed patterns that may be overwhelmed by head ones. These above two innovative designs contribute to considerable improvements together.
@@
Figure 4: Performance comparison over different item groups between RaSeRec and DuoRec. The bar represents NDCG@5, while the line represents the performance improvement percentage of RaSeRec compared with DuoRec.
4.3 Improving Base Backbones (RQ2)
RaSeRec is a model-agnostic retrieval-augmented SeRec paradigm that can enhance various base models. To verify this, instead of using the default Self-attentive backbone SASRec Kang and McAuley (2018), we evaluate RaSeRec using other backbones, i.e., RNN-based GRU4Rec Hidasi et al. (2016) and Bert-based BERT4Rec Sun et al. (2019). As the average performance of DuoRec and MCLRec is close, we mainly compare with DuoRec in the following. As shown in Table 3, RaSeRec performs best with different backbones, indicating that RaSeRec has a good generalization ability. The impressive performance gain demonstrates the effectiveness of our retrieval-augmented mechanism, even though we just add a very small set of parameters (about 5%). In summary, RaSeRec possesses the virtue of the plug-and-play property, which can enhance various given base models with retrieval augmentation while leaving the base model backbone unchanged.
4.4 Long-tailed Recommendation (RQ3)
@@
As stated in Section 1, existing SeRec models make it hard to recall long-tailed patterns via implicit memories due to skewed supervised signal distribution. To verify whether RaSeRec is of promise in solving this issue, we split the test user sequences into 10 groups based on the target item’s popularity measured by the number of interactions, meanwhile keeping the total number of test user sequences in each group the same. And the larger the GroupId, the more popular the target item. The results are shown in Figure 4. From the results, we observe that: (1) Compared with DuoRec, the performance improvements of RaSeRec mostly come from accurately recommending long-tailed and head items. This verifies that retrieval-augmented user representation well recalls both long-tailed and head patterns. (2) We observe that RaSeRec performs relatively worse in recommending general items, i.e., $5$-th and $6$-th groups. We think the main reason is that the relevant memories are not retrieved while the noisy memories are introduced. How to effectively retrieve relevant knowledge is also an open challenge in RAG Zhao et al. (2024b); Wang et al. (2023); Li et al. (2024). In summary, we suggest performing retrieval augmentation when recommending both the long-tailed and head items.
5 Related Works
@@
Retrieval-Augmented Generation. RAG prevails in LLMs, which introduces non-parametric knowledge into LLMs to supplement their incomplete, incorrect, or outdated parametric knowledge Gao et al. (2023). It has been shown highly effective in many research fields, including computer vision Xie et al. (2023); Sharifymoghaddam et al. (2024), speech Xue et al. (2024); Wang et al. (2024a), code Li et al. (2023c); Wang et al. (2024b), and graph Edge et al. (2024). Although being well-studied in many fields, very limited works exploit RAG in SeRec. A very recent one is RUEL Wu et al. (2023). However, it differs from our work in (1) it lacks retrieval training, and (2) it falls short in decoupling implicit and explicit memories, while our work proposes a general collaborative-based and retrieval-augmented training  paradigm for  SeRec.
@@
Sequential Recommendation. Pioneering attempts on SeRec are based on Markov Chain (MC) to model item-item transition relationships, e.g., FPMC Rendle et al. (2010). To capture long-term dependencies, some studies first adopt RNN for sequential recommendations by modeling sequence-level item transitions, e.g., GRU4Rec Hidasi et al. (2016). Simultaneously, CNN has been explored for sequential recommendations by modeling the union-level sequential patterns Tang and Wang (2018). Recently, the success of self-attention models in NLP Vaswani et al. (2017) has spawned a series of Transformer-based sequential recommendation models, e.g., SASRec Kang and McAuley (2018) and BERT4Rec Sun et al. (2019). Considering sparse supervised signals, SSL has been widely adopted in SeRec to alleviate this issue Xie et al. (2022); Qin et al. (2023); Qiu et al. (2022). Enlightened by generative models, some studies have attempted to apply LLMs to solve SeRec Zheng et al. (2024a); Li et al. (2023a, b); Zheng et al. (2024b). Albeit studied for ages, almost all existing methods focus on exploiting implicit memories hidden in the model. This work is in an orthogonal direction. It exploits explicit memories in a retrieval-augmented way and opens a new research direction for SeRec.
6 Conclusion
@@
In this work, we unveiled the limitations of Vanilla and SSL-Augmented SeRec paradigms and explored the potential of RAG to solve the limitations. Particularly, we propose a retrieval-augmented SeRec paradigm, which learns to refine user representation with explicit retrieved memories and adapt to preference drifts by maintaining a dynamic memory bank. Extensive experiments on three datasets manifest the advantage of RaSeRec in terms of recommendation performance, long-tailed recommendation  as well  as  high  customizations.
Limitations
Despite our innovations and improvements, we must acknowledge certain limitations in our work and also reveal some potential directions that warrant in-depth research:
  • Firstly, we perform retrieval training based on ID features, which results in a dataset-centric embedding model that cannot be used in out-of-domain scenarios. In the field of RAG, text-based embedding models can leverage the generality of natural language for common knowledge in different scenarios, which promotes field development to a large extent. As such, we consider it essential to design a generic embedding model for the further development of the  retrieval-augmented  SeRec.
  • Secondly, we treat all training user sequences except the current input user sequence to construct the memory bank $\mathcal{M}$. In Appendix G, we have studied the influence of different partitions of $\mathcal{M}$ on the model performance. The results show that using all parts of $\mathcal{M}$ does not necessarily achieve optimal performance. In RAG, there emerges a lot of work devoted to optimizing data indexing, e.g., the fine-grained chunking strategies Chen et al. (2024). Currently, the data indexing optimization techniques of $\mathcal{M}$ for RaSeRec remain an open field that warrants further investigation.
  • Thirdly, we fixed the number of retrieved memories $K$ during training and inference. In Appendix F, we have studied the impact of different $K$ values on model performance. We find that both too large or small $K$ values hurt the model performance. The main reason is that a large $K$ value may introduce noise, while a small $K$ value may not provide enough information. In RAG, there emerges a lot of work devoted to optimizing retrieval strategies, e.g., retrieval when needed Jiang et al. (2023). Currently, the retrieval techniques for RaSeRec remain an open field that warrants further investigation.
@@ Appendix C Baselines
We compare RaSeRec with the following three groups of competitive baseline models: (1) Non-sequential recommendation models (Non-SeRec):
@@
  • BPR-MF Rendle et al. (2009) uses matrix factorization to model users and items with BPR loss.
@@
  • GRU4Rec Hidasi et al. (2016) uses the Gate Recurrent Unit (GRU) to model user sequence.
  • Caser Tang and Wang (2018) applies Convolutional Neural Networks (CNN) to model union-level sequential patterns for SeRec.
  • SASRec Kang and McAuley (2018) models user sequences via self-attention mechanism.
  • BERT4Rec Sun et al. (2019) applies the bi-directional self-attention mechanism with masked item training to model user sequence.
  • $\rm S^{3}Rec_{MIP}$ Zhou et al. (2020) learns the correlation among items, attributes and etc. Here, we adopt the Mask Item Prediction (MIP) variant.
@@
  • CL4SRec Xie et al. (2022) proposes item crop, mask, and reorder to construct different views of user sequences.
@@
  • CoSeRec Liu et al. (2021) propose two robust data augmentations (i.e., item substitute and insert), to create high-quality views for SSL.
@@
  • ICLRec Chen et al. (2022b) first mines the latent users’ intents and creates SSL signals between sequences and intents.
@@
  • DuoRec Qiu et al. (2022) proposes model-level augmentation with different sets of dropout masks on the model backbone.
@@
  • MCLRec Qin et al. (2023) designs learnable augmenters to generate views and applies meta-learning to guide the training of augmenters.
Figure 5: Parameter sensitivity w.r.t. the control coefficients $\alpha,\beta$ and the number of retrieved memories $K$, where we plot the error bar with the mean $\pm$ $\gamma\times$standard deviation and $\gamma$ is set as 500×500$\times$500 ×the scale interval of the y-axis.
Appendix D Metrics
[…]
    $\begin{split}&\text{NDCG@}N=\frac{1}{|\mathcal{U}|}\sum_{u\in\mathcal{U}}\sum_{j=1}^{N}\\ &\delta(v_{t+1}=R(u)_{j})\frac{1}{\log_{2}(j+1)},\end{split}$ (14)
    where $R(u)_{j}$ denotes the $j$-th recommended item for user $u$.
Configuration Value
number of Transformer layers $L$ 2
number of heads 2
hidden size $d$ 64
maximum sequence length $L$ 50
number of clusters $k$ 128
number of clusters accessed 1
control coefficient $\alpha$* $[0.0,1.0]$
control coefficient $\beta$* $[0.0,1.0]$
number of retrieved memories $K$* $[5,55]$
learning rate 0.001
optimizer Adam
mini-batch size 1024
Table 6: Hyperparameter settings, where * denotes different hyperparameter settings for different datasets.
Appendix E Implementation Details
We use the Adam optimizer Kingma and Ba (2015) to optimize model parameters with the learning rate of 0.001, the mini-batch size of 1024, $\beta_{1}=0.9$, and $\beta_{2}=0.999$, where the maximum number of epochs is set to 100 for both pre-training and fine-tuning stages. We train models with the early stopping strategy that interrupts training if the HR@10 result on the validation set continues to drop until 10 epochs. For RAM, we tune $\alpha$, $\beta$ and $K$ within the ranges of $\{0.0,0.1,0.2,...,1.0\}$, $\{0.0,0.1,0.2,...,1.0\}$ and $\{5,10,15,...,55\}$, respectively, referring to Appendix F for more details. As for Faiss, we set the number of clusters $k$ to 128, set the number of clusters accessed as 1, and adopt the cosine similarity as the measurement to retrieve relevant memories. As for the sequence encoder, we adopt a two-layer Transformer and set the head number as 2 for each self-attention block, where the hidden size $d$ and the maximum sequence length $L$ are set to 64 and 50 respectively. Additionally, for Non-SeRec and Vanilla SeRec methods, we use results reported by DuoRec Qiu et al. (2022) as these methods already have generally acknowledged experimental results on the three benchmark datasets used in this paper. We ran each experiment five times and reported the average. Table 6 summarizes the detailed hyper-parameter settings.
(a) Results on the Beauty dataset w.r.t. HR@5 and NDCG@5.
(b) Results on the Sports dataset w.r.t. HR@5 and NDCG@5.
(c) Results on the Clothing dataset w.r.t. HR@5 and NDCG@5.
Figure 6: Performance comparison when ablating different partitions of the memory bank $\mathcal{M}$. The bar represents HR@5 and NDCG@5, while the line represents the percentage of performance degradation compared to the “All”.
Appendix F Parameter Sensitivity (RQ4)
RQ4: How does the performance of RaSeRec vary with different hyper-parameter values?
Thirdly, we observe that the performance of RaSeRec reaches the peak to different $K$ across three datasets, e.g., 20 is the best value for the Beauty dataset. This demonstrates the effectiveness of augmenting user representation via retrieved explicit memories and manifests that setting a suitable number of retrieved memories can considerably enhance performance. The results on the three datasets follow similar trends, where the performance gradually improves as $K$ increases and then starts to degrade when further increasing $K$. This phenomenon is also observed in long-context RAG Jin et al. (2024); Yue et al. (2024), where the generative modeling performance initially increases and then declines when consistently increasing the number of retrieved passages. The main reason is that increasing the number of retrieved memories will inevitably introduce irrelevant memories (i.e., “noise”) that can mislead the retrieval-augmented module. In conclusion, we suggest tuning $K$ in the range of $10\sim 30$ for the Beauty and Sports datasets, and $30\sim 50$ for the Clothing dataset.
Dataset #len.range #avg.length %proportion
Beauty <3 1.50 34.03%
3-6 4.15 31.18%
>6 18.88 34.79%
Sports <3 1.50 37.56%
3-6 4.14 33.98%
>6 15.31 28.46%
Clothing <2 1.00 24.54%
2-3 2.38 39.38%
>3 8.06 36.08%
Table 7: Statistics of each partition.
(a) Results on the Beauty dataset w.r.t. HR@5 and NDCG@5.
(b) Results on the Sports dataset w.r.t. HR@5 and NDCG@5.
Figure 7: Performance comparison over different user groups between the model with retrieval augmentation and without retrieval augmentation. The bar represents HR@5 and NDCG@5, while the line denotes the performance improvement percentage of “w/ aug” compared to “w/o aug”.
Appendix G Study on Memory Bank (RQ5)
To investigate the influence of different partitions of the memory bank on the model performance, we divide the memory bank $\mathcal{M}$ into three partitions according to the length of user sequences. The statistics of each partition are summarized in Table 7. For convenience, we use $\mathrm{S}$, $\mathrm{M}$, and $\mathrm{L}$ to represent the Short-term, Medium-term, and Long-term memory bank, respectively. For example, on the beauty dataset, we treat user sequences with length less than 3, those between 3 and 6, and those greater than 6 as $\mathrm{S}$, $\mathrm{M}$, and $\mathrm{L}$ respectively, where the proportion of different partitions accounts for about 1/3. The experimental results are shown in Figure 6. “All” denotes using all parts of the memory bank, including $\mathrm{S}$, $\mathrm{M}$, and $\mathrm{L}$, while $\{\mathrm{S},\mathrm{L}\}$ denotes using parts of the memory bank, i.e., $\mathrm{S}$ and $\mathrm{L}$, and so on.
From the experimental results, we observe that different datasets show different preferences toward the combinations of the partitions. Specifically, the beauty dataset prefers “All”; the sports dataset prefers $\{\mathrm{S},\mathrm{M}\}$; the clothing dataset prefers $\{\mathrm{S},\mathrm{L}\}$. We think the main reason is associated with #avg.length. As for the beauty dataset, its average length is longer than the other two datasets and it prefers the $\mathrm{L}$ partition. Specifically, the models with $\mathrm{L}$ partition largely outperform those without $\mathrm{L}$ partition, as shown in Figure 6(a). As for the sports and clothing datasets, their average length is shorter than that of the beauty dataset, and they perform relatively worse when only provided with $\mathrm{L}$ partition, as shown in Figure 6(b) and 6(c). Besides, we observe that the model equipped with parts of the memory bank (e.g., $\{\mathrm{S,M}\}$ is the best for the sports dataset and $\{\mathrm{S,L}\}$ is the best for the clothing dataset) even outperforms that with all parts of the memory bank. This demonstrates that not all memories are beneficial and some ones are harmful to recommendation performance. This is also a challenging issue Cuconasu et al. (2024); Xiang et al. (2024); Fang et al. (2024) in RAG that warrants further research. In a nutshell, we suggest trying different combinations or directly using all of the memories, because “All” usually leads to satisfactory performance.
Appendix H Study on User Interaction Frequency (RQ6)
RQ6: Whether RaSeRec benefits both high-frequency and low-frequency users?

### v3

@@ 4.1.2. Baselines
To verify the effectiveness of RaSeRec, we compare it with the following three groups of competitive models: (1) Non-sequential recommendation models (Non-SeRec) include PopRec and BPR-MF (Rendle et al., 2009); (2) Vanilla sequential recommendation models (Vanilla SeRec) include GRU4Rec (Hidasi et al., 2016), Caser (Tang and Wang, 2018), SASRec (Kang and McAuley, 2018), BERT4Rec (Sun et al., 2019), and $\rm S^{3}Rec_{MIP}$ (Zhou et al., 2020); and (3) SSL-Augmented sequential recommendation models (SSL-Augmented SeRec) include CL4SRec (Xie et al., 2022), CoSeRec (Liu et al., 2021), ICLRec (Chen et al., 2022b), DuoRec (Qiu et al., 2022), and MCLRec (Qin et al., 2023). Detailed descriptions of each baseline model can be seen in Appendix C.
4.1.3. Implementation and Evaluation
@@
We use Recbole (Zhao et al., 2022) to implement baselines. For models with learnable item embedding, we set the hidden size $d=64$. For each baseline, we follow the suggested settings reported in their original paper to set hyper-parameters. More implementation details can be seen in Appendix E. Following the leave-one-out strategy, we hold out the last interacted item of each user sequence for testing, the second last item for validation, and all the earlier ones for training. We measure the recommendation performance via HR and NDCG (see Appendix D), where the cut-off position (@$N$) is set as 5 and 10.
4.2. Comparison with Baselines (RQ1)
@@
Table 2 and 8 summarize the performance of all baselines and RaSeRec w.r.t. @5 and @10, respectively. From the experimental results, we observe that: (1) The performance of the Non-SeRec models is unsurprisingly worse than the vanilla SeRec models, indicating the necessity of modeling sequential patterns for capturing more accurate user preferences. On the other hand, SSL-Augmented SeRec models perform better than those without SSL, indicating the superiority of supplementing the recommendation task with self-supervised learning. (2) Benefiting from the retrieval-augmented mechanism, RaSeRec considerably outperforms all baseline models in 16 out of 18 cases with $p$-value less than 0.005, demonstrating the improvement of RaSeRec over the SOTA baselines is statistically significant. The reasons are mainly twofold: (i) By sustainably maintaining a dynamic memory bank, RaSeRec can quickly adapt to preference drifts. (ii) By retrieving memories, RaSeRec can explicitly recall long-tailed patterns that may be overwhelmed by head ones. These above two innovative designs contribute to considerable improvements together.
4.3. Improving Base Backbones (RQ2)
RaSeRec is a model-agnostic retrieval-augmented SeRec paradigm that can enhance various base models. To verify this, instead of using the default Self-attentive backbone SASRec (Kang and McAuley, 2018), we evaluate RaSeRec using other backbones, i.e., RNN-based GRU4Rec (Hidasi et al., 2016) and Bert-based BERT4Rec (Sun et al., 2019). As the average performance of DuoRec and MCLRec is close, we mainly compare with DuoRec in the following. As shown in Table 3, RaSeRec performs best with different backbones, indicating that RaSeRec has a good generalization ability. The impressive performance gain demonstrates the effectiveness of our retrieval-augmented mechanism, even though we just add a very small set of parameters (about 5%). In summary, RaSeRec possesses the virtue of the plug-and-play property, which can enhance various given base models with retrieval augmentation while leaving the base model backbone unchanged.
@@ 4.3. Improving Base Backbones (RQ2)
Figure 4. Performance comparison over different item groups between RaSeRec and DuoRec. The bar represents NDCG@5, while the line represents the performance improvement percentage of RaSeRec compared with DuoRec.
(a) Results on the Beauty dataset w.r.t. HR@5 and NDCG@5.
(b) Results on the Sports dataset w.r.t. HR@5 and NDCG@5.
(c) Results on the Clothing dataset w.r.t. HR@5 and NDCG@5.
Figure 5. Performance comparison when ablating different partitions of the memory bank $\mathcal{M}$. The bar represents HR@5 or NDCG@5, while the line represents the percentage of performance degradation compared to the “All”.
4.4. Long-tailed Recommendation (RQ3)
@@
As stated in Section 1, existing SeRec models make it hard to recall long-tailed patterns via implicit memories due to skewed supervised signal distribution. To verify whether RaSeRec is of promise in solving this issue, we split the test user sequences into 10 groups based on the target item’s popularity measured by the number of interactions, meanwhile keeping the total number of test user sequences in each group the same. And the larger the GroupId, the more popular the target item. The results are shown in Figure 4. From the results, we observe that: (1) Compared with DuoRec, the performance improvements of RaSeRec mostly come from accurately recommending long-tailed and head items. This verifies that retrieval-augmented user representation well recalls both long-tailed and head patterns. (2) We observe that RaSeRec performs relatively worse in recommending general items, i.e., $5$-th and $6$-th groups. We think the main reason is that the relevant memories are not retrieved while the noisy memories are introduced. How to effectively retrieve relevant knowledge is also an open challenge in RAG (Zhao et al., 2024b; Wang et al., 2023; Li et al., 2024). In summary, we suggest performing retrieval augmentation when recommending the long-tailed and head items.
Table 4. Model performance w.r.t. preference drifts.
Metrics HR NDCG
Datasets Drifts @5 @10 @5 @10
Beauty (A) Full 0.0569 0.0860 0.0369 0.0463
(B) 10% 0.0560 0.0842 0.0360 0.0450
(C) 20% 0.0550 0.0843 0.0355 0.0449
(D) 30% 0.0542 0.0843 0.0350 0.0447
Sports (A) Full 0.0331 0.0497 0.0211 0.0264
(B) 10% 0.0326 0.0492 0.0208 0.0261
(C) 20% 0.0322 0.0490 0.0204 0.0258
(D) 30% 0.0328 0.0490 0.0208 0.0260
4.5. Alleviating Preference Drift (RQ4)
As stated in Section 1, existing SeRec models trained on past data may recommend undesirable items due to preference drift. To verify whether RaSeRec is promising in alleviating this issue, we simulate this scenario by removing a certain ratio (i.e., 10%, 20%, and 30%) of the preference data with the latest timestamp from the memory bank. And, “Full” denotes using all preference data in the memory bank to remedy preference drift. From the results shown in Table 4, we find that (1) In general, the larger the preference drifts, the worse the model performance; (2) “Full” performs best as it can remedy preference drift by maintaining a dynamic memory bank with the latest preference data; and (3) In a few cases, large preference drift is not necessarily worse performance, e.g., (D) on the sports dataset. The main reason is different datasets have different preferences towards short-, medium-, and long-term memories, which also greatly affects performance (more details can be seen in §4.6). In a nutshell, the above observations verify that RaSeRec can alleviate preference drift effectively.
Table 5. Statistics of each partition.
Dataset #len.range #avg.length %proportion
Beauty ¡3 1.50 34.03%
3-6 4.15 31.18%
¿6 18.88 34.79%
Sports ¡3 1.50 37.56%
3-6 4.14 33.98%
¿6 15.31 28.46%
Clothing ¡2 1.00 24.54%
2-3 2.38 39.38%
¿3 8.06 36.08%
4.6. Study on Memory Bank (RQ5)
To investigate the influence of different partitions of the memory bank on the model performance, we divide the memory bank $\mathcal{M}$ into three partitions according to the length of user sequences. The statistics of each partition are summarized in Table 5. For convenience, we use $\mathrm{S}$, $\mathrm{M}$, and $\mathrm{L}$ to represent the Short-term, Medium-term, and Long-term memory bank, respectively. For example, on the beauty dataset, we treat user sequences with length less than 3, those between 3 and 6, and those greater than 6 as $\mathrm{S}$, $\mathrm{M}$, and $\mathrm{L}$ memory bank respectively, where the proportion of different partitions accounts for about 1/3. The experimental results are shown in Figure 5. “All” denotes using all parts of the memory bank, including $\mathrm{S}$, $\mathrm{M}$, and $\mathrm{L}$, while $\{\mathrm{S},\mathrm{L}\}$ denotes using parts of the memory bank, i.e., $\mathrm{S}$ and $\mathrm{L}$, and so on.
(a) Results on Beauty w.r.t. HR@5 and NDCG@5.
(b) Results on Sports w.r.t. HR@5 and NDCG@5.
Figure 6. Model performance w.r.t. noise ratio on Beauty and Sports datasets. The bar represents HR@5 or NDCG@5, while the line represents the percentage of performance degradation compared to the test user sequences without added noise.
From the experimental results, we observe that different datasets show different preferences toward the combinations of the partitions. Specifically, the beauty dataset prefers “All”; the sports dataset prefers $\{\mathrm{S},\mathrm{M}\}$; the clothing dataset prefers $\{\mathrm{S},\mathrm{L}\}$. We think the main reason is associated with #avg.length.
As for the beauty dataset, its average length is longer than the other two datasets and it prefers the $\mathrm{L}$ partition. Specifically, the models with $\mathrm{L}$ partition largely outperform those without $\mathrm{L}$ partition, as shown in Figure 5(a). As for the sports and clothing datasets, their average length is shorter than that of the beauty dataset, and they perform relatively worse when only provided with $\mathrm{L}$ partition, as shown in Figure 5(b) and 5(c). Besides, we observe that the model equipped with parts of the memory bank (e.g., $\{\mathrm{S,M}\}$ is the best for the sports dataset and $\{\mathrm{S,L}\}$ is the best for the clothing dataset) even outperforms that with all parts of the memory bank. The above observations demonstrate that not all memories are beneficial and some ones are harmful to recommendation performance. This is also a challenging issue (Cuconasu et al., 2024; Xiang et al., 2024; Fang et al., 2024) in RAG that warrants further research. In a nutshell, we suggest trying different combinations or directly using all of the memories, because “All” usually leads to satisfactory performance. In this work, we use “All” by default.
4.7. Robustness against Noise Data (RQ6)
In real-world scenarios, recommender systems usually suffer from the noisy interaction issue (Gao et al., 2022; Tian et al., 2022), e.g., a user clicked an item by mistake. As such, it is necessary to verify RaSeRec’s robustness against noisy data. To simulate this scenario, we train models with full training data and randomly add a certain ratio (i.e., 10%, 20%, and 30%) of negative items into each test user sequence. The experimental results are shown in Figure 6, where models with retrieval augmentation mention “w/ aug” otherwise “w/o aug”. From the experimental results, we observe that adding noises significantly degrades the performance of models with or without retrieval augmentation. However, “w/ aug” consistently outperforms “w/o aug” under any noise ratio in terms of HR@5 and NDCG@5. Moreover, the blue line is always below the green line, as shown in Figure 6, demonstrating that “w/ aug” can endow the model backbone with more robustness against the noisy interactions during inference. For example, on the Beauty and Sports datasets, “w/ aug” with a 20% noise ratio outperforms “w/o aug” with a 10% noise ratio. Moreover, on the Sports dataset, “w/ aug” drops 14.7% of its original performance, while “w/o aug” drops 18.4%, in terms of NDCG@5, when both have a 10% noise ratio. This fully implies that with retrieval augmentation, RaSeRec can retrieve useful sequential patterns as references to alleviate the influence of noisy interactions and make a suitable recommendation. In a nutshell, we suggest performing retrieval augmentation to alleviate the negative influence of noisy interactions.
5. Related Works
5.1. Retrieval-Augmented Generation
@@
Recently, RAG prevails in LLMs, introducing non-parametric knowledge into LLMs to supplement their incomplete, incorrect, or outdated parametric knowledge (Gao et al., 2023). It has been shown highly effective in many research fields, not limited to natural language, including computer vision (Xie et al., 2023; Sharifymoghaddam et al., 2024; Rao et al., 2022, 2023), speech (Xue et al., 2024; Wang et al., 2024b), code (Li et al., 2023a; Wang et al., 2024a), and graph (Edge et al., 2024). Although being well-studied in many fields, very limited works exploit RAG in recommender systems, let alone SeRec. A very recent one is RUEL (Wu et al., 2023). However, it differs from our work in (1) it lacks retrieval training, and (2) it falls short in decoupling implicit and explicit memories, while our work proposes a model-agnostic collaborative-based pre-training and retrieval-augmented fine-tuning paradigm for SeRec that can be applied to any ID-based SeRec backbones and improve their performance.
5.2. Sequential Recommendation
@@
Pioneering attempts on SeRec are based on Markov Chain (MC) to model item-item transition relationships, e.g., FPMC (Rendle et al., 2010). To capture long-term dependencies, some studies first adopt RNN for SeRec by modeling sequence-level item transitions, e.g., GRU4Rec (Hidasi et al., 2016). Simultaneously, CNN has been explored for SeRec by modeling the union-level sequential patterns (Tang and Wang, 2018). Recently, the success of self-attention models in NLP (Vaswani et al., 2017) has spawned a series of Transformer-based sequential recommendation models, e.g., SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019). Considering sparse supervised signals, SSL has been widely adopted in SeRec to alleviate this issue (Xie et al., 2022; Qin et al., 2023; Qiu et al., 2022). Enlightened by generative models, some studies have attempted to apply LLMs to solve SeRec (Zheng et al., 2024b; Li et al., 2023b, c; Zheng et al., 2024a). Albeit studied for ages, almost all existing methods focus on exploiting implicit memories hidden in the model. This work is in an orthogonal direction. It exploits explicit memories in a retrieval-augmented manner and opens a new research direction for SeRec.
6. Conclusion and Future Works
@@
In this work, we unveiled the limitations of the Vanilla SeRec and SSL-Augmented SeRec paradigms and explored the potential of RAG to solve them. In particular, we propose a retrieval-augmented SeRec paradigm, which learns to refine user representation with explicit retrieved memories and adapt to preference drifts by maintaining a dynamic memory bank. Extensive experiments on three datasets manifest the advantage of RaSeRec in terms of recommendation performance, long-tailed recommendation, and alleviating preference drift. Despite our innovations and improvements, we recognize some issues that warrant further study. In this work, we fixed the number of retrieved memories $K$ during training and inference. In Appendix F, we have studied the impact of different $K$ values on the model performance. We find both too large and small $K$ values hurt the model performance. Furthermore, in Section 4.4 and Appendix G, we have studied the impact of retrieval augmentation on different item groups and user groups. The experimental results show that retrieval augmentation may hurt the performance of some groups. These observations motivate us to develop an active retrieval augmentation mechanism for RaSeRec, i.e., retrieval when needed.
@@ Appendix C Baselines
We compare RaSeRec with the following three groups of competitive baseline models:
(1) Non-sequential recommendation models (Non-SeRec):
@@
  • BPR-MF (Rendle et al., 2009) uses matrix factorization to model users and items with BPR loss.
@@
  • GRU4Rec (Hidasi et al., 2016) uses the Gate Recurrent Unit (GRU) to model user sequence.
  • Caser (Tang and Wang, 2018) applies Convolutional Neural Networks (CNN) to model union-level sequential patterns for SeRec.
  • SASRec (Kang and McAuley, 2018) models user sequences via self-attention mechanism.
  • BERT4Rec (Sun et al., 2019) applies the bi-directional self-attention mechanism with masked item training to model user sequence.
  • $\rm S^{3}Rec_{MIP}$ (Zhou et al., 2020) learns the correlation among items, attributes and etc. Here, we adopt the Mask Item Prediction (MIP) variant.
@@
  • CL4SRec (Xie et al., 2022) proposes item crop, mask, and reorder to construct different views of user sequences.
@@
  • CoSeRec (Liu et al., 2021) propose two robust data augmentations (i.e., item substitute and insert), to create high-quality views for SSL.
@@
  • ICLRec (Chen et al., 2022b) first mines the latent users’ intents and creates SSL signals between sequences and intents.
@@
  • DuoRec (Qiu et al., 2022) proposes model-level augmentation with different sets of dropout masks on the model backbone.
@@
  • MCLRec (Qin et al., 2023) designs learnable augmenters to generate views and applies meta-learning to guide the training of augmenters.
Table 8. Performance comparison (@10), where the best results are boldfaced and the second-best ones are underlined.